water is exposed to infrared radiation of wavelengths 2.8x 10^-4 cm. assume that all the radiation is absorbed and converted into heat. how many protons are required to raise the temperature of 2.0 g of water by 2.0 K.
1
Expert's answer
2017-12-16T04:25:07-0500
λ = 2.8* 10-4 cm = 2.8* 10^­-6 m
m= 2*10-3kg
ΔT = 2 K
c= 4200 J/kg*K
cl= 3*10^8m/s
h=6.62*10-34 J*s
Q=c*m* ΔT = N*E
E=h*ν = h* cl/ λ
N= c*m* ΔT* λ/h* cl
N=(4200 J/kg*K *2*10^-3kg*2 K*2.8* 10^­-6 m)/ 6.62*10^-34 J*s* 3*10^8m/s = 2.37 * 10^20